# Predicting the Performance of Black-box LLMs through Self-Queries

### v2

@@ 5 Discussion
Overall, we find that this work better enables the safer usage of black-box LLMs, as we can audit their performance with predictors (that even have performance guarantees) trained on the representations extracted via QueRE. For instance, an interesting line of future work that can apply QueRE in data curation pipelines, as a means of getting lightweight predictors of synthetic data quality. We also remark that we believe our work provides insight into the explainability (or lack thereof) of black-box LLMs. Extracting representations by asking a model questions eliciting is, in some sense, an evaluation of its ability to meaningfully understand its own behavior and respond to prompts. However, we note that this is an imperfect comparison, as these extracted features are treated in an uninterpretable manner (i.e., as features for a supervised learning model). In fact, our finding that responses to random sequences can reveal information about model behavior aligns with prior work describing flaws in existing interpretability frameworks (Friedman et al., 2023; Singh et al., 2024). We believe that our work provides promising results towards the deployment of LLMs as components of larger systems or in providing supervision and highlighting new insights into the understanding of these (black-box) LLMs.
@@ A.1 Full Table Results
We present the full set of our results on open-ended QA tasks (Table 4) and closed-ended QA tasks (Table 5) comparing all different methods on all LLMs applied to all considered datasets.
@@
Dataset LLM Logits RepE Pre-conf Post-conf Answer P. QueRE
NQ LLaMA3-3B 0.5933 0.6639 0.5265 0.8186 0.9070 0.9596
LLaMA3-8B 0.5626 0.6521 0.5148 0.8502 0.7923 0.9483
LLaMA3-70B 0.6663 0.7124 0.5563 0.7976 0.8328 0.9527
GPT-3.5 0.5700* - 0.5429 0.6025 0.5088 0.6714
GPT-4o-mini 0.5463* - 0.5395 0.6060 0.5033 0.6654
SQuAD LLaMA3-3B 0.6893 0.7033 0.5081 0.9220 0.6856 0.9579
LLaMA3-8B 0.6843 0.6993 0.5145 0.7928 0.6061 0.9492
LLaMA3-70B 0.6983 0.7068 0.5099 0.7818 0.7532 0.8944
GPT-3.5 0.5597* - 0.5074 0.5822 0.4990 0.6685
GPT-4o-mini 0.6468* - 0.5015 0.5078 0.5753 0.7092
Table 5: AUROC in predicting model performance on closed-ended QA tasks. “-” denotes unreported values or that RepE cannot be applied to black-box models; “*” denotes that Full Logits for GPT-3.5 is a sparse vector with nonzero values only for the top-5 logits.
Dataset LLM Logits RepE Pre-conf Post-conf Answer P. Sem U. QueRE
@@
A.2 Uncertainty Quantification Baselines
Another line of work in uncertainty quantification [Xiong et al., 2023] looks to extract estimates of model confidence from the LLM directly. This is fundamentally related to our problem setting, but perhaps is less focused on the applications of predicting model behavior (and certainly not focused on our other applications of detecting adversarial models or distinguishing between architectures). These baselines include: (1) Vanilla confidence elicitation, which is to directly ask the model for a confidence score, (2) TopK, asking the LLM for its TopK answer options with their corresponding confidences, (3) CoT, asking the LLM to first explain its reasoning step-by-step before asking for a confidence score, and (4) Multistep, which asks the LLM to produce multiple steps of reasoning each with a confidence score. We use $K=3$ for the TopK baseline and 3 steps in the multistep baseline.
@@ A.2 Uncertainty Quantification Baselines
We observe that QueRE achieves stronger performance than these these uncertainty quantification baselines (Table 6). We also remark that QueRE is more widely applicable as these methods (which are implemented in Xiong et al. [2023]), as they heavily on being able to parse the format of responses for closed-ended QA tasks. On the contrary, QueRE indeed applies to open-ended QA tasks (see our strong results in Figure 2).
A.3 Studying the Role of Diversity in Elicitation Questions
We also provide experiments to study the exact role of diversity in these elicitation questions, on top of our prior experiment using random sequences. We use various prompts to generate other types of elicitation questions (see Section D.3 for the resulting questions). One prompt attempts to produce a set of more diverse elicitation questions, while another attempts to output a set of more similar elicitation questions.
@@ A.3 Studying the Role of Diversity in Elicitation Questions
We analyze the performance of these approaches in generating elicitation questions that differ in human interpretable notions of diversity (Figure 9). We observe that generally, attempting to increase diversity does not necessarily improve performance. This suggests that as it is difficult for us to interpret what diversity is important for these LLMs, and that the notion of diversity generated through prompting for more “diverse” questions does not necessarily result in diverse features extracted from the LLM. We believe that better understanding this discrepancy in notions of “diversity” is an interesting line for future research.
A.4 Extended Random Sequence Ablations
Building upon our previous ablation on using random sequences in our approach instead of elicitation questions, we vary the number of random sequences and elicitation questions to better understand the impact and importance of diversity in the follow-up questions/prompts to the model.
Figure 10: Comparison of using varying amounts of prompts of random sequences of natural language or elicitation questions in QueRE. The results are presented on the LLaMA3-8B model from left-to-right as: Squad, NQ, and HaluEval.
We observe that using elicitation questions generally achieves better performance (Figure 10). We also note that after a certain number of elicitation prompts, increasin the number of random sequences can lead to more benefits when compared to elicitation questions, likely due to the greater diversity in their construction. These experiments reveal that at a certain budget of API queries / number of elicitation prompts, the elicitation questions are more efficient at achieving strong performance when compared to random sequences of natural text. However, random sequences of natural text have more diversity and can later match (or potentially even exceeded) the performance of elictation questions given a sufficient number of follow-up sequences. This reveals that indeed natural sequences of language can extract useful information from these models in a black-box manner, which we believe is an interesting result. This also suggests that generating prompts for QueRE is extremely easy, as they can take on the form of unrelated sequences of language.
A.5 Additional Generalization Results
For our PAC-Bayes bounds over linear models [Jiang et al., 2019], we use a prior over weights of $\mathcal{N}(0,\sigma^{2}I)$, giving us our bound as
$E\left[L(\beta)\right]\leq E\left[\hat{L}(\beta)\right]+\sqrt{\frac{\frac{||w||_{2}^{2}}{4\sigma^{2}}+\log\frac{n}{\delta}+10}{n-1}}$
@@ A.5 Additional Generalization Results
We remark that our work defines a different line to approach generalization bounds through a more human-interactive approach to eliciting low-dimensional representations, although we remark that this human-interaction in specifying these elicitation questions must be independent of any training data (e.g., questions must be predefined before seeing the dataset of interested). Perhaps the most related work in this line are existing works that have achieved tight generalization bounds for VLMs [Akinwande et al., 2023] and for LLMs modeling log-likelihoods [Lotfi et al., 2023].
A.6 Additional Calibration Results
We present the ECE comparison between QueRE and the Answer Probs baseline on the remaining datasets (Figure 11). We observe similar behavior, in that QueRE defines much more calibrated predictors than simply using Answer Probs in almost every case.
A.7 Robustness to System Prompts
We provide an additional experiment to illustrate that QueRE is robust to slight changes in the system prompt. We have two sets of vectors extracted via QueRE from a GPT-4o-mini model without an additional system prompt, and a version with an additional system prompt that is ”You are a helpful and cautious assistant.” on the Boolean Questions dataset.
@@ A.7 Robustness to System Prompts
A.8 Representation Visualizations by Different Modle Sizes
We also provide visualizations of our extracted embeddings for various LLMs architectures, noting that different models are distinctly clustered in the plots (Figure 12).
@@ A.8 Representation Visualizations by Different Modle Sizes
A.9 Results using MLPs
We provide experiments that use 5-layer MLPs instead of linear classifiers to predict model performance, where each of the MLP hidden layers are of size 8. We compare different methods that extract representations (that are not single dimensional). We observe that performance is still stronger with QueRE, showing that the benefits still hold for models other than linear classifiers (Table 8).
@@ A.9 Results using MLPs
A.10 Additional Results for Varying the Number of Elicitation Questions
We present additional results when varying the number of elicitation questions on other QA tasks. Here, we only look at subsets of the elicitation questions and do not include the components of preconf, postconf and answer probabilities. We observe that across all tasks, we observe a consistent increase in performance as we increase the size of the subset of elicitation questions that we consider, with diminishing benefits as we have a larger number of prompts (Figure 13). Generally, increasing the number of elicitation prompts leads to an increase in AUROC, clearly defining a tradeoff between extracting the most informative black-box representation and the overall cost of introducing more queries to the LLM API. An interesting future question is how to best select elicitation questions, and perhaps, removing those that add redundant information or noise. This is reminiscent of work in prior work in pruning or weighting ensembles of weak learners [Mazzetto et al., 2021a, b] or in dimensionality reduction [Van Der Maaten et al., 2009].
@@ Appendix B Proof of Proposition 1
We again present Proposition 1 and now include its proof in its entirety.
@@
Consider the standard logistic regression setup (as in the work of Stefanski and Carroll [1985]), where we are learning a linear model $\beta$, which satisfies that
@@
In our setting, instead of having access to covariates $X_{i}$, we rather have access to an approximation of these covariates $\hat{X}_{i}$, which is an average of $k$ samples from Ber($X_{i}$). An application of the results in the work of Stefanski and Carroll [1985] gives us the result that the MLE $\hat{\beta}$ is a consistent estimator of $\beta_{0}$, given that $k\to\infty$. This is fairly straightforward as when $k\to\infty$, we have that $\frac{1}{k}\sum_{j=1}^{k}\hat{X}_{i}^{j}\to X_{i}$, implying that the noise in the covariates goes to 0 as $n\to\infty$ (i.e., satisfying a main condition of the result in Stefanski and Carroll [1985]).
@@
We use this term to analyze $||(\beta-\beta_{0})||_{2}$. First, we can apply the Cauchy-Schwarz inequality, which gives us that
$||\beta-\beta_{0}||_{2}\leq\left|\left|\frac{\partial\beta}{\partial X}\right|\right|_{F}\cdot||\hat{X}-X||_{2},$
Then, we note that $||\hat{X}-X||_{2}$ converges to 0 at a rate of $O\left(\sqrt{\frac{d}{k}}\right)$ via an application of the CLT. We can also analyze the term
$\left|\left|\frac{\partial\beta}{\partial X}\right|\right|_{F}$ $\leq\left|\left|(X^{T}DX)^{-1}\right|\right|_{F}\cdot\left|\left|\frac{1}{n}\Delta X^{T}(\sigma-y)+X^{T}D\Delta X\beta\right|\right|_{F}$
@@
$\left|\left|(X^{T}DX)^{-1}\right|\right|_{F}\leq\frac{\sqrt{d}}{\sigma_{min}(X^{T}DX)}$
@@
$\left|\left|\frac{1}{n}\Delta X^{T}(\sigma-y)\right|\right|_{F}=\sqrt{\frac{d}{k}}$
@@
Now, assuming that $\beta$ has norm $||\beta||^{2}\leq B$, we have that
@@
$\left|\left|\frac{\partial\beta}{\partial X}\right|\right|_{F}$ $\leq\left(\frac{\sqrt{d}}{\sigma_{min}(X^{T}DX)}\right)\left(\sqrt{\frac{d}{k}}+\sqrt{\frac{Bd}{n}}\right)$
@@
when we note that $d$ is roughly a constant with respect to $n,k$, and $B$ is a constant, and assuming that $\sigma_{min}(X^{T}DX)=O(\frac{1}{\sqrt{n}})$. Putting this back together with the Taylor expansion and the standard asymptotics of $||\hat{X}-X||$, we get that $\beta$ converges to $\beta_{0}$ at a rate of $O\left(\frac{\sqrt{n}}{k}\right)$.

### v4

@@ 5 Discussion
Our contributions find that querying a language model with follow-up questions leads to features that are useful in a wide variety of applications in predicting model behavior. Remarkably, they can often match the performance of predictors that work in the white-box setting over model internals when predicting correctness on LLM benchmarks or in detecting when language models have been adversarially manipulated. Overall, we believe that our work provides promising results towards reliably predicting the behavior of language models and detecting when they have been adversarially manipulated, which supports the potential of their deployment in larger autonomous systems and foundations towards more trustworthy language models.
Limitations
While QueRE demonstrates strong predictive performance across many tasks, it has a few limitations. First, although the features extracted via QueRE are grounded in natural language, our focus is not on interpretability or attribution. We treat these features purely as abstract inputs to a predictive model, rather than as explanations or understanding of model behavior. Second, our approach introduces latency through multiple follow-up queries per example, although this can be mitigated through batching. Finally, while our method generalizes across datasets and model families, it relies on the assumption that a model’s responses to follow-up questions meaningfully vary—a property that may not hold for very low-quality language models.
Acknowledgements
DS is supported by the National Science Foundation Graduate Research Fellowship Program under Grant No DGE2140739. The authors would also like to thank Yewon Byun for their thoughtful feedback and helpful suggestions on the introduction and problem setting.
@@ A.1 Full Table Results
We present the full set of our results on open-ended QA tasks (Table˜5) and closed-ended QA tasks (Table˜6) comparing all different methods on all LLMs applied to all considered datasets.
@@
Dataset LLM Logits RepE Pre-conf Post-conf Self-Consis. Sem. Entropy QueRE
NQ LLaMA3-3B 0.5933 0.6639 0.5265 0.8186 0.6245 0.6659 0.9596
LLaMA3-8B 0.5626 0.6521 0.5148 0.8502 0.5314 0.6327 0.9483
LLaMA3-70B 0.6663 0.7124 0.5563 0.7976 0.6291 0.6661 0.9527
GPT-3.5 0.6567* - 0.5941 0.6693 0.6695 0.7063 0.6755
GPT-4o-mini 0.5459* - 0.6277 0.6778 0.6956 0.6880 0.6780
SQuAD LLaMA3-3B 0.6893 0.7033 0.5081 0.9220 0.5714 0.5192 0.9579
LLaMA3-8B 0.6843 0.6993 0.5145 0.7928 0.5343 0.5207 0.9492
LLaMA3-70B 0.6983 0.7068 0.5099 0.7818 0.5280 0.5014 0.8944
GPT-3.5 0.6173* - 0.5061 0.5392 0.6639 0.5290 0.6899
GPT-4o-mini 0.7413* - 0.5043 0.5899 0.7203 0.5246 0.7113
Table 6: AUROC in predicting model performance on closed-ended QA tasks. “-” denotes unreported values or that RepE cannot be applied to black-box models; “*” denotes that Full Logits for GPT-3.5 is a sparse vector with nonzero values only for the top-5 logits. We bold the best performing black-box method, and italicize the best white-box method when it outperforms the black-box approaches.
Dataset LLM Logits RepE Pre-conf Post-conf Answer P. Sem. Entropy QueRE
@@
A.2 Distinguishing Between Levels of Quantization
We also provide additional experiments that can distinguish between different levels of quantization of a language model. We find that on SQuAD with LLaMA models, we find that QueRE can easily distinguish between model responses that are generated via different levels of quantization, while these other baselines fail.
Table 7: Comparison of different quantization settings (4-bit, 16-bit) against full 32-bit precision for LLaMA3-3B and LLaMA3-8B models. We report AUROC scores for various uncertainty and representation-based detectors. QueRE remains consistently strong across quantization settings.
Model (Quantization) Pre-conf Post-conf Logprobs Semantic Ent. QueRE
LLaMA3-3B (4bit vs 32bit) 0.61 0.59 0.57 0.71 0.99
LLaMA3-3B (16bit vs 32bit) 0.50 0.51 0.51 0.58 0.99
LLaMA3-8B (4bit vs 32bit) 0.67 0.50 0.61 0.63 0.98
LLaMA3-8B (16bit vs 32bit) 0.51 0.54 0.55 0.56 0.97
A.3 Uncertainty Quantification Baselines
Another line of work in uncertainty quantification [xiong2023can] looks to extract estimates of model confidence from the LLM directly. This is fundamentally related to our problem setting, but perhaps is less focused on the applications of predicting model behavior (and certainly not focused on our other applications of detecting adversarial models or distinguishing between architectures). These baselines include: (1) Vanilla confidence elicitation, which is to directly ask the model for a confidence score, (2) TopK, asking the LLM for its TopK answer options with their corresponding confidences, (3) CoT, asking the LLM to first explain its reasoning step-by-step before asking for a confidence score, and (4) Multistep, which asks the LLM to produce multiple steps of reasoning each with a confidence score. We use $K=3$ for the TopK baseline and 3 steps in the multistep baseline.
@@ A.3 Uncertainty Quantification Baselines
We observe that QueRE achieves stronger performance than these these uncertainty quantification baselines (Table˜8). We also remark that QueRE is more widely applicable as these methods (which are implemented in xiong2023can), as they heavily on being able to parse the format of responses for closed-ended QA tasks. On the contrary, QueRE indeed applies to open-ended QA tasks (see our strong results in Figure 2).
A.4 Models Trained on QueRE are Better Calibrated
While we have previously reported the AUROC of our predictors, we are also interested in the calibration of our models (e.g., accuracy at a given confidence threshold). This is particularly useful for high-stakes settings, when we may only want to defer prediction to a LLM when we are confident in its performance. We observe that predictors defined by QueRE generally have much lower ECE compared to those defined by using answer probabilities.
Our approach shows promise in constructing well-calibrated and performant predictors of LLM performance, which are important for the application of LLMs in high-stakes settings [weissler2021role, thirunavukarasu2023large].
A.5 Studying the Role of Diversity in Follow-up Questions
We also provide experiments to study the exact role of diversity in these elicitation questions, on top of our prior experiment using random sequences. We use various prompts to generate other types of follow-up questions (see Section˜D.3 for the resulting questions). One prompt attempts to produce a set of more diverse queries, while another attempts to output a set of more similar queries.
@@ A.5 Studying the Role of Diversity in Follow-up Questions
We analyze the performance of these approaches in generating elicitation questions that differ in human interpretable notions of diversity (Figure˜7). We observe that generally, attempting to increase diversity does not necessarily improve performance. This suggests that as it is difficult for us to interpret what diversity is important for these LLMs, and that the notion of diversity generated through prompting for more “diverse" questions does not necessarily result in diverse features extracted from the LLM. We believe that better understanding this discrepancy in notions of “diversity” is an interesting line for future research.
A.6 Unrelated Sequences Ablations
We also explore the potential of, instead of using follow-up questions, to use unrelated sequences of natural langauge. We vary the number of these unrelated sequences of language and elicitation questions to better understand the impact and importance of diversity in the follow-up questions/prompts to the model.
Figure 8: Comparison of using varying amounts of prompts of unrelated sequences of natural language or follow-up questions in QueRE. The results are presented on the LLaMA3-8B model from left-to-right as: Squad, NQ, and HaluEval.
We observe that using follow-up questions generally achieves better performance (Figure˜8). However, we still find that indeed unrelated sequences of language can extract useful information from these models in a black-box manner, which we believe is an interesting result. This suggests that generating prompts for QueRE is extremely easy, as they can take on the form of unrelated sequences of language and do not need to be limited to the form or follow-up questions. In fact, our finding that responses to unrelated sequences can reveal information about model behavior aligns with prior work describing flaws in existing interpretability frameworks [friedman2023interpretability, singh2024rethinking].
A.7 Additional Results for Distinguishing Models
Figure 9: Accuracy in distinguishing representations from LLMs of different sizes on SQuAD.
We now present additional results on distinguishing between different model sizes on the SQuAD dataset. We observe the same trends, finding that QueRE better distinguishes between different LLaMA3 and GPT models, when compared to alternatives.
A.8 Additional Generalization Results
For our PAC-Bayes bounds over linear models [jiang2019fantastic], we use a prior over weights of $\mathcal{N}(0,\sigma^{2}I)$, giving us our bound as
[…]
@@ A.8 Additional Generalization Results
We remark that our work defines a different line to approach generalization bounds through a more human-interactive approach to eliciting low-dimensional representations, although we remark that this human-interaction in specifying these elicitation questions must be independent of any training data (e.g., questions must be predefined before seeing the dataset of interested). Perhaps the most related work in this line are existing works that have achieved tight generalization bounds for VLMs [akinwande2023understanding] and for LLMs modeling log-likelihoods [lotfi2023non].
A.9 Robustness to System Prompts
We provide an additional experiment to illustrate that QueRE is robust to slight changes in the system prompt. We have two sets of vectors extracted via QueRE from a GPT-4o-mini model without an additional system prompt, and a version with an additional system prompt that is "You are a helpful and cautious assistant.” on the Boolean Questions dataset.
@@ A.9 Robustness to System Prompts
A.10 Representation Visualizations by Different Model Sizes
We also provide visualizations of our extracted embeddings for various LLMs architectures, noting that different models are distinctly clustered in the plots (Figure˜10).
@@ A.10 Representation Visualizations by Different Model Sizes
A.11 Results Using MLPs
We provide experiments that use 5-layer MLPs instead of linear classifiers to predict model performance, where each of the MLP hidden layers are of size 8. We compare different methods that extract representations (that are not single dimensional). We observe that performance is still stronger with QueRE, showing that the benefits still hold for models other than linear classifiers (Table˜10).
@@ A.11 Results Using MLPs
A.12 Additional Results for Varying the Number of Elicitation Questions
We present additional results when varying the number of elicitation questions on other QA tasks. Here, we only look at subsets of the elicitation questions and do not include the components of preconf, postconf and answer probabilities. We observe that across all tasks, we observe a consistent increase in performance as we increase the size of the subset of follow-up questions that we consider, with diminishing benefits as we have a larger number of prompts (Figure˜11). Generally, increasing the number of elicitation prompts leads to an increase in AUROC, clearly defining a tradeoff between extracting the most informative black-box representation and the overall cost of introducing more queries to the LLM API. An interesting future question is how to best select follow-up queries, and perhaps, removing those that add redundant information or noise. This is reminiscent of work in prior work in pruning or weighting ensembles of weak learners [mazzetto2021adversarial, mazzetto2021semi] or in dimensionality reduction [van2009dimensionality].
@@ A.12 Additional Results for Varying the Number of Elicitation Questions
A.13 Latency Analysis
We additionally report latency–performance trade-offs for QueRE on SQuAD with LLaMA3-8B, varying the number of follow-up questions up to the full 50 used in our experiments. Table 11 compares QueRE against key black-box baselines, including Post-Conf, Self-Consistency, and Semantic Entropy.
Table 11: Latency–performance trade-offs for LLaMA3-8B on SQuAD. We report AUROC and average runtime per example (in seconds).
Method AUROC Avg. Runtime (s)
Post-Conf 0.515 0.08
QueRE (5 follow-ups) 0.868 0.17
QueRE (10 follow-ups) 0.897 0.17
QueRE (20 follow-ups) 0.916 0.36
QueRE (30 follow-ups) 0.928 0.55
QueRE (40 follow-ups) 0.933 0.74
QueRE (50 follow-ups) 0.949 0.89
Self-Consistency 0.534 0.19
Semantic Entropy 0.521 2.44
We find that QueRE consistently outperforms other approaches at similar or lower runtimes, demonstrating a superior latency–accuracy trade-off. For a comparable latency to Self-Consistency (0.17 s vs. 0.19 s), QueRE with just 5–10 follow-up questions achieves dramatically higher AUROC ($\sim$0.90 vs. 0.53). Furthermore, QueRE significantly outperforms the much slower Semantic Entropy baseline, achieving both higher predictive power and greater computational efficiency.
A.14 Precision and F1 on Incorrect Examples
We additionally compute precision and F1 scores on negative samples (i.e., cases where the LLM produces incorrect answers). Higher precision on these examples indicates more reliable detection of incorrect model behavior. Table 12 reports results for LLaMA3-8B and LLaMA3-70B across all QA benchmarks.
Table 12: Precision and F1 on negative (incorrect) examples across datasets. Each cell reports precision / F1.
Dataset LLM Pre-Conf Post-Conf Answer P. Sem. Entropy QueRE
BoolQ LLaMA3-8B 0.324 / 0.435 0.307 / 0.440 0.266 / 0.317 0.334 / 0.462 0.446 / 0.569
LLaMA3-70B 0.410 / 0.509 0.410 / 0.529 0.325 / 0.362 0.427 / 0.550 0.591 / 0.684
CS QA LLaMA3-8B 0.804 / 0.567 0.821 / 0.553 0.857 / 0.609 0.970 / 0.485 0.920 / 0.836
LLaMA3-70B 0.807 / 0.765 0.843 / 0.710 0.817 / 0.735 0.891 / 0.928 0.953 / 0.943
HaluEval LLaMA3-8B 0.741 / 0.551 0.711 / 0.635 0.761 / 0.723 0.712 / 0.817 0.803 / 0.798
LLaMA3-70B 0.772 / 0.632 0.775 / 0.787 0.787 / 0.680 0.794 / 0.675 0.810 / 0.800
DHate LLaMA3-8B 0.374 / 0.411 0.508 / 0.540 0.373 / 0.516 0.444 / 0.546 0.747 / 0.761
LLaMA3-70B 0.394 / 0.380 0.456 / 0.437 0.360 / 0.506 0.491 / 0.433 0.785 / 0.777
Across nearly all datasets and both model scales, QueRE remains the strongest black-box method, achieving the highest precision and F1 on negative examples. This further confirms that QueRE offers a more reliable mechanism for identifying incorrect or uncertain model behavior, complementing its superior AUROC performance.
A.15 Multi-Negative Identification
We further extend our evaluation to a multi-negative identification setting, where the goal is to determine whether a given response originates from GPT-4o-mini, with negatives drawn from a pool of three other models (LLaMA3-8B, LLaMA3-70B, and GPT-3.5). This setup more closely reflects real-world scenarios such as detecting fraudulent API substitutions or model impersonation. As shown in Table 13, QueRE remains highly effective, achieving near-perfect accuracy while maintaining strong generalization across datasets.
Table 13: Multi-negative identification results. The task is to identify whether a response is from GPT-4o-mini among negatives from LLaMA3-8B, LLaMA3-70B, and GPT-3.5. We report AUROC.
Dataset Pre-Conf Post-Conf Answer Probs Sem. Entropy QueRE
SQuAD 0.541 0.522 0.501 0.580 0.998
BoolQ 0.501 0.532 0.564 0.592 0.998
Across both benchmarks, QueRE achieves substantially higher AUROC than all black-box baselines, highlighting its robustness in distinguishing target model generations even in the presence of multiple distractor models. These findings underscore QueRE’s practical utility for tasks that require reliable model source identification and detection of potentially substituted or spoofed model outputs.
@@ Appendix B Proof of Proposition 1
We again present Proposition˜1 and now include its proof in its entirety.
@@
Consider the standard logistic regression setup (as in the work of stefanski1985covariate), where we are learning a linear model $\beta$, which satisfies that
@@
In our setting, instead of having access to covariates $X_{i}$, we rather have access to an approximation of these covariates $\hat{X}_{i}$, which is an average of $k$ samples from Ber($X_{i}$). An application of the results in the work of stefanski1985covariate gives us the result that the MLE $\hat{\beta}$ is a consistent estimator of $\beta_{0}$, given that $k\to\infty$. This is fairly straightforward as when $k\to\infty$, we have that $\frac{1}{k}\sum_{j=1}^{k}\hat{X}_{i}^{j}\to X_{i}$, implying that the noise in the covariates goes to 0 as $n\to\infty$ (i.e., satisfying a main condition of the result in stefanski1985covariate).
@@
We use this term to analyze $(\beta-\beta_{0})||_{2}$. First, we can apply the Cauchy-Schwarz inequality, which gives us that
[…]
Then, we note that $\hat{X}-X||_{2}$ converges to 0 at a rate of $O\left(\sqrt{\frac{d}{k}}\right)$ via an application of the CLT. We can also analyze the term
[…]
@@
[…]
@@
[…]
@@
Now, assuming that $\beta$ has norm $\beta||^{2}\leq B$, we have that
@@
[…]
@@
when we note that $d$ is roughly a constant with respect to $n,k$, and $B$ is a constant, and assuming that $\sigma_{min}(X^{T}DX)=O(\frac{1}{\sqrt{n}})$. Putting this back together with the Taylor expansion and the standard asymptotics of $\hat{X}-X||$, we get that $\beta$ converges to $\beta_{0}$ at a rate of $O\left(\frac{\sqrt{n}}{k}\right)$.